A sum of Rs 50000 amounts to Rs 75000 at a certain rate of simple interest in 5 yrs. Find the sum of money that amounts to Rs 96800 at the same rate of compound interest in 2 yrs.

Grade:10

1 Answers

Arun
4 years ago
Dear student
 
25000 = 50000* r/100 * 5
 
r = 10 %
 
now 96800 = x( 1 + 10/100)^2
 
96800/x = 1.21
x = 80000
 
Hope it helps